OpenStudy (anonymous):

parallel to y=-2x+1 thorugh the point (0,3)

OpenStudy (wolfe8):

Well when 2 lines are parallel, what can be said about their gradients? How can you obtain the gradient from a straight-line equation? Lastly, there is a formula for constructing line equations for when you have the gradient and a point the line passes through.

OpenStudy (anonymous):

hint: find the slope of the first equation by looking at \[y = mx + b\] use that slope, along with the given points, and use the point slope form \[(y-y _{1}) = m(x-x _{1})\]solve for x.